Stock XYZ has a current dividend of $7.00 . The dividend is expected to grow forever at a rate of 1.80% . Based on the riskiness of XYZ, its discount rate is 2.00% With this information, what is the dividend yield from today to period 1?
Answer #1

Current Dividend = D0 = $ 7, Growth Rate = g = 1.8 % and Discount Rate = r = 2 %

Expected Dividend = D1 = 7 x 1.018 = $7.126

Current price = P0 = 7.126 / (r-g) = 7.126 / (0.02 - 0.018) = $ 3563

Dividend Yield = (D1/P0) = (7.126/3563) = 0.002 or 0.2 %